Suppose the following demand and supply function:

Qd = 750 – 25P

Qs = -300 + 20 P


      i.         Find equilibrium price and quantity

    ii.         Find consumer and producer surplus



1
Expert's answer
2021-05-27T18:57:31-0400

i. At equilibrium:

"Q_s=Q_s \\

-300 +20P=750 -25P \\

20P+25P = 750 +300 \\

45P = 1050 \\

P_{eq} = \frac{1050}{45}=23.33 \\

Q_{eq} = -300 +20 \times 23.33 \\

= -300 + 466.66 \\

= 166.66"

ii.

"Q_d = 750 -25P"

If "Q_d = 0, P_d = 30"

Consumer surplus "= \frac{1}{2}(P_d-P_{eq})Q_{eq}"

"= \frac{1}{2} \times(30-23.33)166.66 = 555.611 \\

Q_s = -300 + 20P"

If "Q_s = 0, P_s=15"

Producer surplus "= \frac{1}{2}(P_{eq} -P_s)Q_{eq}"

"= \frac{1}{2}(23.33-15)166.66 = 693.889"